OpenStudy (kity_kisses1234):

Pleaseeee help me .... The half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass. Starting with 160 grams of a radioactive isotope, how much will be left after 4 half-lives? round your answer to the nearest gram.

OpenStudy (mrnood):

same method 1 year = 160 * 0,5 2 year 160 *0.5^2 4 years...?

OpenStudy (kity_kisses1234):

10 grams?

OpenStudy (mrnood):

sorry - slight mistake above: same method 1 half-life= 160 * 0,5 2 half-lives =160 *0.5^2 4 half-lives...?

OpenStudy (mrnood):

yes = 0.5 ^4 = 1/16 so 160/16 = 10 g